# GLBench: A Comprehensive Benchmark for Graph with Large Language Models

### v1

@@ 1 Introduction
Graphs are ubiquitous in modeling the relational and structural aspects of real-world objects, encompassing a wide range of real-world scenarios  [18]. Many of these graphs have nodes that are associated with text attributes, resulting in the emergence of text-attributed graphs, such as citation graphs  [15, 32] and web links  [33]. For example, in citation graphs, each node represents a paper, and its corresponding textual description (e.g., title and abstract) is treated as the node’s text attributes.
To tackle graphs with both node attributes and graph structural information, conventional pipelines typically fall into two categories. Firstly, graph neural networks (GNNs)  [23, 43, 10] have emerged as the dominant approach through recursive message passing and aggregation mechanisms among nodes. They often utilize non-contextualized shallow embeddings, such as bag-of-words  [11] and skip-gram  [34], as shown in the previous benchmarks  [38, 15]. Secondly, pretrained language models (PLMs) can be directly employed to encode the text associated with each node, transforming the problem into a text classification task without considering graph structures  [52, 7]. Following  [47], we refer to these two learning paradigms as GNN-based and PLM-based methods.
In recent years, the advent of large language models (LLMs) with massive context-aware knowledge and semantic comprehension capabilities (e.g., LLaMA  [42], GPTs  [1], Mistral  [19]) marks a significant advancement in AI research, also leading a notable shift in the way we interact with graphs. For example, LLMs can retrieve semantic knowledge that is relevant to the nodes to improve the quality of initial node embeddings of GNNs  [13, 45]. LLMs can also encode graphs through carefully designed prompts and directly make predictions in an autoregressive manner  [5, 16]. In addition, LLMs can be aligned with GNNs in the same vector space, enabling GNNs to be more semantically aware  [52, 29]. In this paper, we refer to this new paradigm of using LLMs to assist with graph-related problems as GraphLLM. We follow existing surveys  [24, 20] to organize existing GraphLLM methods into three categories based on the role (i.e., enhancer, predictor, and aligner) played by LLMs throughout the entire model pipeline.
@@ 2.3 LLMs for Graphs
Role Method Predictor GNN PLM/LLM Techniques Used Learning Scenarios Venue Code Fine-tune Prompt Supervised Zero-shot Enhancer GIANT  [9] GNN GraphSAGE, etc. BERT ICLR’22 Link TAPE  [13] GNN RevGAT ChatGPT ICLR’24 Link OFA  [26] GNN R-GCN Sentence-BERT ICLR’24 Link ENGINE  [54] GNN GraphSAGE LLaMA-2 IJCAI’24 Link ZeroG  [25] GNN SGC Sentence-BERT SIGKDD’24 Link Predictor InstructGLM  [50] LLM - FLAN-T5/LLaMA-v1 EACL’24 Link GraphText  [53] LLM - ChatGPT/GPT-4 Arxiv Link GraphAdapter  [17] LLM GraphSAGE LLaMA-2 WWW’24 Link GraphGPT  [40] LLM GT Vicuna SIGIR’24 Link LLaGA  [6] LLM - Vicuna/LLaMA-2 ICML’24 Link Aligner GLEM  [52] GNN/LLM GraphSAGE, etc. RoBERTa ICLR’23 Link Patton  [21] LLM GT BERT/SciBERT ACL’23 Link
@@ 3.2 Research Questions
Experiment Design. To answer this research question, we evaluate three GraphLLM methods with zero-shot capabilities on all datasets in GLBench for node classification, which involves an optional pre-training phase using source datasets that have non-overlapping classes with the target datasets. To ensure data distribution practices in NLP and CV, where source datasets are typically larger than target datasets, we select the two largest graphs in GLBench, i.e., Ogbn-arxiv and Reddit for pre-training, and use the remaining datasets for inference. We involve additional models as baselines including (i) graph self-supervised learning (SSL), which focus on the transferability of learned structures, such as DGI  [44] and GraphMAE  [14]; (ii) LLMs, which only consider semantic information for classification, such as LLaMA3-70B  [42], GPT-3.5-turbo  [35], GPT-4o  [1], and DeepSeek-chat  [3]. We follow  [7] to design input prompts for each dataset; (iii) Emb w/ NA, a simple training-free baseline that utilizes both structural and semantic information for zero-shot inference.
@@ 4.1 Supervised Performance Comparison (RQ1)
There is no obvious scaling law in existing GraphLLM methods. Recently, neural scaling law on graphs has become a focal point of discussion  [4, 28], particularly in terms of two perspectives: data scaling and model scaling. Since we ensure consistency in the data, we only focus on the model scaling law. However, from Table  4, it can be observed that the performance does not clearly scale with model size. For instance, within the LLM-as-enhancer style, ENGINE  [54] employs 7B-scale LLMs, yet it does not demonstrate a significant performance gap compared to the use of 110M-scale LLMs like OFA  [26]. It is more noticeable in LLM-as-predictor methods, with models like InstructGLM  [50] and GraphText  [53] using 7B-scale LLMs yield unsatisfactory results, indicating that merely accumulating parameters does not guarantee better performance.
@@ 5 Conclusion
Acknowledgement. We want to thank Haitao Mao at Michigan State University, Xinni Zhang at The Chinese University of Hong Kong, Wenxuan Huang at Zhejiang University, Ruosong Ye at Rutgers University, and Caiqi Zhang at Cambridge University for their valuable insights into the field of GraphLLM, and constructive comments on this paper.
@@ B.1 Hyperparameter Tuning
  • For ENGINE, we follow the instruction from the original paper  [54], use
@@
  • For GraphText, we follow the instruction from the original paper  [53], use
@@
  • For GraphAdapter, we follow the instruction from the original paper  [17], use
@@
  • For GraphGPT, we follow the instruction from the original paper  [40], use
@@ C.2 Zero-shot Inference
Appendix D Broader Discussion
D.1 Limitations of GLBench

### v3

@@ 1 Introduction
Graphs are ubiquitous in modeling the relational and structural aspects of real-world objects, encompassing a wide range of real-world scenarios  [57, 11, 59, 58, 49, 64]. Many of these graphs have nodes that are associated with text attributes, resulting in the emergence of text-attributed graphs, such as citation graphs  [17, 35] and web links  [36]. For example, in citation graphs, each node represents a paper, and its textual description (e.g., title and abstract) is treated as the node’s text attributes.
To tackle graphs with both node attributes and graph structural information, conventional pipelines typically fall into two categories. Firstly, graph neural networks (GNNs)  [24, 12, 27, 56] have emerged as the dominant approach through recursive message passing and aggregation mechanisms among nodes. They often utilize non-contextualized shallow embeddings, such as bag-of-words  [13] and skip-gram  [37], as shown in the previous benchmarks  [41, 17]. Secondly, pretrained language models (PLMs) can be directly employed to encode the text associated with each node, transforming the problem into a text classification task without considering graph structures  [60, 8]. Following  [51], we refer to these two learning paradigms as GNN-based and PLM-based methods.
In recent years, the advent of large language models (LLMs) with massive context-aware knowledge and semantic comprehension capabilities (e.g., LLaMA  [45], GPTs  [1], Mistral  [20]) marks a significant advancement in AI research, also leading a notable shift in the way we interact with graphs  [7]. For example, LLMs can retrieve semantic knowledge that is relevant to the nodes to improve the quality of initial node embeddings of GNNs  [15, 48]. LLMs can also encode graphs through carefully designed prompts and directly make predictions in an autoregressive manner  [5, 18]. In addition, LLMs can be aligned with GNNs in the same vector space, enabling GNNs to be more semantically aware  [60, 32]. In this paper, we refer to this new paradigm of using LLMs to assist with graph-related problems as GraphLLM. We follow existing surveys  [25, 21] to organize existing GraphLLM methods into three categories based on the role (i.e., enhancer, predictor, and aligner) played by LLMs throughout the entire model pipeline.
@@ 2.3 LLMs for Graphs
Role Method Predictor GNN PLM/LLM Techniques Used Learning Scenarios Venue Code Fine-tune Prompt Supervised Zero-shot Enhancer GIANT  [10] GNN GraphSAGE, etc. BERT ICLR’22 Link TAPE  [15] GNN RevGAT ChatGPT ICLR’24 Link OFA  [29] GNN R-GCN Sentence-BERT ICLR’24 Link ENGINE  [63] GNN GraphSAGE LLaMA-2 IJCAI’24 Link ZeroG  [26] GNN SGC Sentence-BERT KDD’24 Link Predictor InstructGLM  [54] LLM - FLAN-T5/LLaMA-v1 EACL’24 Link GraphText  [61] LLM - ChatGPT/GPT-4 Arxiv Link GraphAdapter  [19] LLM GraphSAGE LLaMA-2 WWW’24 Link GraphGPT  [43] LLM GT Vicuna SIGIR’24 Link LLaGA  [6] LLM - Vicuna/LLaMA-2 ICML’24 Link Aligner GLEM  [60] GNN/LLM GraphSAGE, etc. RoBERTa ICLR’23 Link Patton  [22] LLM GT BERT/SciBERT ACL’23 Link
@@ 3.2 Research Questions
Experiment Design. To answer this research question, we evaluate three GraphLLM methods with zero-shot capabilities on all datasets in GLBench for node classification, which involves an optional pre-training phase using source datasets that have non-overlapping classes with the target datasets. To ensure data distribution practices in NLP and CV, where source datasets are typically larger than target datasets, we select the two largest graphs in GLBench, i.e., Ogbn-arxiv and Reddit for pre-training, and use the remaining datasets for inference. We involve additional models as baselines including (i) graph self-supervised learning (SSL), which focus on the transferability of learned structures  [28], such as DGI  [47] and GraphMAE  [16]; (ii) LLMs, which only consider semantic information for classification, such as LLaMA3-70B  [45], GPT-3.5-turbo  [38], GPT-4o  [1], and DeepSeek-chat  [3]. We follow  [8] to design input prompts for each dataset; (iii) Emb w/ NA, a simple training-free baseline that utilizes both structural and semantic information for zero-shot inference.
@@ 4.1 Supervised Performance Comparison (RQ1)
There is no obvious scaling law in existing GraphLLM methods. Recently, neural scaling law on graphs has become a focal point of discussion  [4, 31], particularly in terms of two perspectives: data scaling and model scaling. Since we ensure consistency in the data, we only focus on the model scaling law. However, from Table  4, it can be observed that the performance does not clearly scale with model size. For instance, within the LLM-as-enhancer style, ENGINE  [63] employs 7B-scale LLMs, yet it does not demonstrate a significant performance gap compared to the use of 110M-scale LLMs like OFA  [29]. It is more noticeable in LLM-as-predictor methods, with models like InstructGLM  [54] and GraphText  [61] using 7B-scale LLMs yield unsatisfactory results, indicating that merely accumulating parameters does not guarantee better performance. In addition, we design a more rigorous and accurate experiment to better analyze the scaling law, as detailed in Appendix  D. This experiment is based on different scales of GNNs and LLMs. The results indicate that it is challenging to identify clear scaling laws, whether related to the scaling of GNNs or LLMs, which is consistent with the insights presented here.
@@ 5 Conclusion
Acknowledgement. This work was supported by National Key Research and Development Program of China Grant No. 2023YFF0725100 and Guangzhou-HKUST(GZ) Joint Funding Scheme 2023A03J0673. We want to thank Haitao Mao at Michigan State University, Xinni Zhang and Zhixun Li at The Chinese University of Hong Kong, Wenxuan Huang at Zhejiang University, Ruosong Ye at Rutgers University, and Caiqi Zhang at Cambridge University for their valuable insights into the field of GraphLLM, and constructive comments on this paper.
@@ B.1 Hyperparameter Tuning
  • For ENGINE, we follow the instructions from the original paper  [63], use
@@
  • For GraphText, we follow the instructions from the original paper  [61], use
@@
  • For GraphAdapter, we follow the instructions from the original paper  [19], use
@@
  • For GraphGPT, we follow the instructions from the original paper  [43], use
@@ C.2 Zero-shot Inference
Appendix D Discussion of Scaling Law
According to  [31], the neural scaling law on graphs “describes how model performance grows with the scale of training, which have two basic forms: data scaling law (how the performance changes with training dataset size) and model scaling law (how the performance changes with model size)”. Since we maintain the dataset splitting in a fair comparison benchmark, we only discuss the model scaling law here. To be specific, we select OFA  [29] and GLEM  [60], which perform well across all datasets in GLBench, for our scaling law experiments. Since these models typically comprise both a GNN and an LLM, we explore scaling law separately for each component. To ensure a fair comparison, we adopt consistent experimental setups across different scales of GNNs and LLMs, such as “learning_rate” and “batch_size” (more details on parameters can be found in Appendix  B.1), with the only variable being the scales of GNNs and LLMs.
Table 9: OFA and GLEM results on Cora dataset with different layers and parameter sizes of GNNs.
Models Layers Metric Units per Layer 128 256 512 768 1024 OFA  [29] 4-layer #Param. 592769 2168577 8269313 18302209 32267265 Acc. 66.10 69.15 72.05 71.13 72.97 5-layer #Param. 691457 2562561 9843713 21843457 38561793 Acc. 64.31 68.67 72.44 74.52 75.00 6-layer #Param. 790145 2956545 11418113 25384705 44856321 Acc. 64.17 69.54 75.10 73.45 74.42 GLEM  [60] 2-layer #Param. 66966165 67562517 68755221 69947925 71140629 Acc. (GNN) 80.95 79.01 81.48 81.24 80.51 Acc. (LLM) 72.68 73.66 72.87 72.05 72.53 3-layer #Param. 67114773 68154645 71119125 75263253 80587029 Acc. (GNN) 80.66 81.24 81.29 81.78 81.43 Acc. (LLM) 71.47 72.92 71.71 73.07 73.07 4-layer #Param. 67263381 68746773 73483029 80578581 90033429 Acc. (GNN) 81.77 82.01 82.16 81.53 82.06 Acc. (LLM) 72.73 71.47 71.95 72.01 72.20
D.1 Different scales of GNNs
We utilize fixed LLMs (i.e., SentenceBERT) and explore the scaling laws of GNNs in GraphLLM methods using GNNs with different layers and dimensions of hidden states. Detailed results for OFA and GLEM are shown in Table  9 and Figure  4, respectively. From the results, we observe that as the number of parameters increases, whether due to an increase in layers or the dimensions, there is no significant trend of improvement in model performance.
Figure 4: Model scaling behaviors of OFA and GLEM with varying model depths. $R^{2}$ close to 1 indicates that the model has a strong explanatory power. $R^{2}$ close to 0 means that the model explains or predicts 0% of the relationship between the dependent and independent variables.
Table 10: OFA and GLEM results on Cora dataset with different parameter sizes of LLMs.
TinyBERT BERT-base BERT-large Sent-BERT_small Sent-BERT RoBERTa_base RoBERTa_large LLaMA2-7B LLaMA2-13B Method 4M 109M 335M 23M 66M 125M 355M 7B 13B GNN LLM GNN LLM GNN LLM GNN LLM GNN LLM GNN LLM GNN LLM GNN LLM GNN LLM OFA 56.33 - 65.52 - 65.18 - 78.97 - 75.24 - 56.00 - 58.03 - 78.72 - 80.56 - GLEM 72.15 37.52 81.57 73.11 79.06 70.60 78.00 69.05 81.23 72.92 76.74 70.89 79.26 72.53 OOM OOM OOM OOM
D.2 Different scales of LLMs
We use fixed GNNs and explore the scaling laws of LLMs in GraphLLM methods using LLMs ranging from 4M to 13B. The results are shown in Table  10. For fairness, we introduce several versions of the same LLM-series for comparison, such as BERT-base and BERT-large. It can be observed that increasing the scale of LLMs has a less noticeable impact on model performance. Although OFA with LLaMA2-13B achieves the best results, the performance does not increase steadily with the increase in parameters, e.g., OFA with RoBERTa-large only achieves 58.3% Accuracy.
In summary, it is hard to identify clear scaling laws, whether related to the scaling of GNNs or LLMs, which is consistent with the insights in GLBench.
Appendix E Broader Discussion
E.1 Limitations of GLBench
@@ E.2 Potential Impacts
Checklist
  1. 1.
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes]
    2. (b)
      Did you describe the limitations of your work? [Yes]
    3. (c)
      Did you discuss any potential negative societal impacts of your work? [Yes]
    4. (d)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes]
  2. 2.
    If you are including theoretical results…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? [N/A]
    2. (b)
      Did you include complete proofs of all theoretical results? [N/A]
  3. 3.
    If you ran experiments (e.g. for benchmarks)…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes]
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes]
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [No]
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes]
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
    1. (a)
      If your work uses existing assets, did you cite the creators? [Yes]
    2. (b)
      Did you mention the license of the assets? [Yes]
    3. (c)
      Did you include any new assets either in the supplemental material or as a URL? [Yes]
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [Yes]
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [Yes]
  5. 5.
    If you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots, if applicable? [N/A]
    2. (b)
      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A]
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [N/A]